An equation and two ordered pairs are given show that each pair is a solution of the equation. Then graph the two pairs to determine another solution.
2y+x=5;(-1,3), (7,-1)

Answer by mananth(16949)

2y+x=5;(-1,3), (7,-1)
2y+x=5;(-1,3),
2*3+(-1)=5
6-1=5
5=5
...
2y+x=5 (7,-1)
2*(-1)+7=5
-2+7=5
5=5
...